OpenStudy (nadiahossain):

The area of the Southern ocean is about 7.85*10^6 square miles. The difference between the areas of the Indian Ocean and the Southern Ocean is about 1.865*10^7 square miles. Explain how to find the area of the Indian Ocean. Then find the area.

Vocaloid (vocaloid):

first rewrite one of the two areas such that: the exponents are the same so 7.85 * 10^6 = 0.785 * 10^7 then just subtract 1.865 - 0.785 = ?

OpenStudy (nadiahossain):

Multiply or Divide?

Vocaloid (vocaloid):

uh, no, just subtract

Vocaloid (vocaloid):

1.865 - 0.785 = ?

OpenStudy (nadiahossain):

But they are asking for the difference

OpenStudy (nadiahossain):

Why the equal sign?

Vocaloid (vocaloid):

difference means subtract

OpenStudy (nadiahossain):

Yeah I know but, 7.85 * 10^6 = 0.785 * 10^7 why is their an equal sign?

Vocaloid (vocaloid):

because the two sides are equal?

OpenStudy (nadiahossain):

And they're asking for area, don't you have to multiply?

Vocaloid (vocaloid):

you would only need to multiply if they gave you width and length, and wanted to find area they already give you the area, so you subtract to find the difference

OpenStudy (nadiahossain):

So to find the area you need to subtract?

Vocaloid (vocaloid):

"The difference between the areas of the Indian Ocean and the Southern Ocean is about 1.865*10^7 square miles. " means Indian Ocean - Southern Ocean = 1.865*10^7, with me so far?

OpenStudy (anonymous):

But wait Vocaloid, the Indian ocean is bigger so wouldn't it be \[\large \sf Indian~ocean-Southern~ocean=1.865 \times 10^{7}\]\[\large \sf Indian~ocean=(1.865+.785) \times 10^{7}\]

OpenStudy (anonymous):

I'm too slow with latex /:

OpenStudy (nadiahossain):

So what will be the explanation for finding the area?

OpenStudy (anonymous):

Well if the difference is.785 * 10^7, then if you add that onto the smaller area, you will get the area of the larger area, which is the Indian Ocean.

OpenStudy (nadiahossain):

So to find the area you need to subtract Indian ocean and Southern Ocean...

OpenStudy (anonymous):

No no no, you have 3 values total, 2 that you know and 1 that you don't. The ones you know are the Southern Ocean's area (the small area) and the difference in area between the small and large areas (.785 * 10^7). So to find the Indian ocean's area (the large area) you will add the difference to the southern ocean's area

OpenStudy (nadiahossain):

And the area would be 1.08*10^7?

OpenStudy (nadiahossain):

wait, how did you get .785? it was 7.85

OpenStudy (anonymous):

That area of the indian ocean would be (1.865+.785) * 10^7 . We got .785 because we changed the 10^6 to 10^7 which is multiplying it by 10, so we had to divide the 7.85 by 10 as well, bringing it down to .785.

OpenStudy (nadiahossain):

so the area is 2.65*10^7 ?

OpenStudy (anonymous):

yes, that is the area of the indian ocean.

OpenStudy (nadiahossain):

oh okay, but I am sorry if I am getting you mad but can you clarify something to me a little more please

OpenStudy (anonymous):

No problem, I'm not mad (not at you at least lol). What is it?

OpenStudy (nadiahossain):

So how is it .785 because if you do this (7.85*10^6)-(1.865*10^7) and you have to add one with the 6 which will be 7 but you have to move more forward...

OpenStudy (nadiahossain):

I think I am wrong somewhere, correct me please

OpenStudy (anonymous):

Well first, it would be a + between the parentheses. Second, when we had 7.85 * 10^6, we changed the numbers on both sides, by multiplying 10^6 by 10 and dividing 7.85 by 10. This changed it to .785 * 10^7 but it kept the ratio the same. Once we had both sides with 10^7 as a constant, we could just add 1.865 and .785, which would be multiplied by the constant 10^7.

OpenStudy (anonymous):

I hope that explained it well enough cause I gtg. have a good night/day/evening etc.

OpenStudy (nadiahossain):

Okay thanks for the help

OpenStudy (anonymous):

ye no problem (: